OpenStudy (anonymous):

Identify the correct x-intercept and y-intercept of the graph of the equation 15x - 5y = 30. intercept = (-2, 0); y-intercept = (0, 6) intercept = (0, -6); y-intercept = (2, 0) x-intercept = (2, 0); y-intercept = (0, -6) intercept = (0, 6); y-intercept = (-2, 0)

OpenStudy (anonymous):

this equation can be rearranged.. 15x - 30 = 5y so y = 3x - 6 y intercept = -6 from here slope = 3 so the third option will be correect

OpenStudy (ghazi):

to find x intercept put y=0 in the equation 15x-5y=30 and to find y intercept put x=0